# Avoiding Catastrophe in Online Learning by Asking for Help

### v5

@@ 2 Related work
Multiplicative objectives. Although online learning traditionally studies the sum of payoffs, there is some work which aims to maximize the product of payoffs (or equivalently, the sum of logarithms). See, e.g., Chapter 9 of Cesa-Bianchi & Lugosi (2006). However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds. Also, like most online learning work, those results assume that payoffs are observed on every time step. In contrast, our agent only receives feedback in response to queries (Table 1) and never observes payoffs. Barman et al. studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product. Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero. This distinction is closely related to the difference between subconstant and sublinear regret.
@@ 3 Model
Inputs. Let $\mathbb{N}$ denote the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs. In the fully adversarial setting, each $x_{t}$ can have arbitrary (possibly randomized) dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $\mathcal{D}_{t}$ from which $x_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $S\subseteq\mathcal{X}$, $\mathcal{D}(S)\leq\frac{1}{\sigma}U(S)$. (In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution; see Definition 1 of Block et al. (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions and payoffs. Let $\mathcal{Y}$ be a finite set of actions. There also exists a special action $\tilde{y}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff. Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$. We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions. Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$. Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive. Unless otherwise noted, all expectations are over any randomization in the agent’s decisions, any randomization in $\boldsymbol{x}$, and any randomization in the adaptive choice of $\boldsymbol{\mu}$.
@@
Local generalization. We assume that $\boldsymbol{\mu}$ and $\pi^{m}$ satisfy local generalization. Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$. Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. This represents the ability to transfer knowledge between similar inputs:
$\big{|}\!\!\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\!\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}\!\!\!\!\!\!\!\big{|}\,\leq\>\underbrace{L||x-x^{\prime}||}_{\text{Input similarity}}$
@@
when all payoffs are strictly positive. To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu^{m}(x_{t})\geq\mu_{0}^{m}$ always. Thus only the agent’s payoffs can be zero, so we can safely define $R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\infty$ whenever $\mu_{t}(x_{t},y_{t})=0$ for some $t\in[T]$. We write $R_{T}^{\times}=R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
The assumption of $\mu^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal. In fact, we argue that high-stakes AI applications should employ a mentor who is almost always safe, i.e., $\mu_{0}^{m}\approx 1$. If no such mentor exists for some application, perhaps that application should be avoided altogether.
@@
Misc. The diameter of a set $S\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(S)=\max_{x,x^{\prime}\in S}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
input $x$payoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
@@ 4.2 Formal definition of construction
$\mu_{f,\boldsymbol{a}}(x,y)=\begin{cases}1&\textnormal{ if }y=a_{j(x)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(x)}-x|\right)&\textnormal{ if }y\neq a_{j(x)}\\ \end{cases}$
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions. This section presents a simple algorithm which guarantees subconstant regret (both multiplicative and additive) and sublinear queries under either of those assumptions. Formal proofs are deferred to Appendix B but we provide intuition and a proof sketch here.
@@ 5.2 Local generalization
Local generalization is vital: this is what allows us to detect when an input is unfamiliar. Crucially, our algorithm does not need to know how inputs are encoded in $\mathbb{R}^{n}$ and does not need to know $L$: it only needs to be able to compute the nearest-neighbor distance $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||$. Thus we only need to assume that there exists some encoding satisfying local generalization.
@@
     $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$  (see Definition 5.4)
@@
[…]
@@ 6 Conclusion and future work
Remaining technical questions. First, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension). Second, the time complexity of Algorithm 1 currently stands at a hefty $\Omega(|\tilde{\Pi}|)$ per time step plus the time to compute $\tilde{\Pi}$. In the standard online learning setting, Block et al. (2022) and Haghtalab et al. (2022) show how to replace discretization approaches like ours with oracle-efficient approaches, where a small number of calls to an optimization oracle are made per round. We are optimistic about leveraging such techniques to obtain efficient algorithms in our setting.
@@ A.3 Proof
Recall the standard inequalities $11-\frac{1}{a}\leq\log a\leq a-$ 1 for any $a>0$.
@@
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
[…]
$=$ $\ R_{T}^{\times}$ (Definition of $(\text{Definition of $R_{T}^{\times}$})$
@@
$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of RT× given $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
[…]
$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
@@
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $-\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$. Also, $W/2\leq E[Z]/2$ implies $\Pr[Z\leq W/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
@@
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
@@
[…]
@@
[…]
@@
    Their “actions” correspond to policies in our setting, not actions in $\mathcal{Y}$. Their number of actions $n$ corresponds to $|\tilde{\Pi}|$.
@@
  function HedgeWithQueries$(p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1])$
@@
     for $t$ from $1$ to $T$ do
[…]
@@ B.2 Main proof
$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\tilde{\pi})]\leq$ $\ \frac{\log|\tilde{\Pi}|}{p^{2}}$
$=$ $\ T\varepsilon\log|\tilde{\Pi}|$
@@
Consider an arbitrary $t\in[T]$. If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ trivially, so assume $t\not\in Q_{T}$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$. Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $||a-b||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
@@
Next, fix a $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the multiset of observed inputs whose mentor action is $y$. Also let $\hat{X}_{2}=\{x_{t}:t\in\hat{Q}_{2}\}$ be the multiset of inputs associated with time steps in $\hat{Q}_{2}$. Note that $|\hat{X}_{2}|=|\hat{Q}_{2}|$, since $\hat{X}_{2}$ is a multiset. We claim that $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$. Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried. Then $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$. Also, since $x,x^{\prime}\in\hat{X}_{2}$ we have $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$. Therefore
$\min_{(x^{\prime\prime},y^{\prime\prime})\in S_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
By Lemma B.5, there exists a ball $B_{1}$ of radius $R:=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.4 implies that
@@
(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound. Therefore
$\operatorname*{\mathbb{E}}[|Q_{T}|]\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
@@
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$

### v6

@@ 2 Related work
Multiplicative objectives. Although online learning traditionally studies the sum of payoffs, there is some work which aims to maximize the product of payoffs (or equivalently, the sum of logarithms). See, e.g., Chapter 9 of Cesa-Bianchi & Lugosi (2006). However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds. Also, like most online learning work, those results assume that payoffs are observed on every time step. In contrast, our agent only receives feedback in response to queries (Table 1) and never observes payoffs. Barman et al. (2023) studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product. Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero. This distinction is closely related to the difference between subconstant and sublinear regret.
@@ 3 Model
Inputs. Let $\mathbb{N}$ denote the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs. In the fully adversarial setting, each $x_{t}$ can have arbitrary (possibly randomized) dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $\mathcal{D}_{t}$ from which $x_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $S\subseteq\mathcal{X}$, $\mathcal{D}(S)\leq\frac{1}{\sigma}U(S)$. (In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution UU. [^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution; see Definition 1 of Block et al. (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions and payoffs. Let $\mathcal{Y}$ be a finite set of actions. There also exists a special action $\tilde{y}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff. Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$. We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions. Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$. Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive. Unless otherwise noted, all expectations are over any randomization in the agent’s decisions, any randomization in $\boldsymbol{x}$, and any randomization in the adaptive choice of $\boldsymbol{\mu}$.
@@
[…]
@@
when all payoffs are strictly positive. To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu_{t}^{m}(x_{t})\geq\mu_{0}^{m}$ always. Thus only the agent’s payoffs can be zero, so we can safely define $R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\infty$ whenever $\mu_{t}(x_{t},y_{t})=0$ for some $t\in[T]$. We write $R_{T}^{\times}=R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
The assumption of $\mu_{t}^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal. In fact, we argue that high-stakes AI applications should employ a mentor who is almost always safe, i.e., $\mu_{0}^{m}\approx 1$. If no such mentor exists for some application, perhaps that application should be avoided altogether.
@@
[…]
input xxpayoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}\dots$1011
@@ 4.2 Formal definition of construction
[…]
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension dd. [^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions. This section presents a simple algorithm which guarantees subconstant regret (both multiplicative and additive) and sublinear queries under either of those assumptions. Formal proofs are deferred to Appendix B but we provide intuition and a proof sketch here.
@@ 5.2 Local generalization
[…]
@@
  $\tilde{\Pi}\leftarrow$ any adversarial cover of $\Pi$ of size at most $(eT/d)^{d}$  (see Definition 5.4)
@@
  with probability $p:$ $\texttt{hedgeQuery}\leftarrow\texttt{true}$
  with probability 1p:1-p: $\texttt{hedgeQuery}\leftarrow\texttt{false}$
@@ 6 Conclusion and future work
[…]
@@ A.3 Proof
Recall the standard inequalities $1-\frac{1}{a}\leq\log a\leq a-1$ for any $a>0$.
@@
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Definition of $R_{T}^{+}$})$
$=$ $\ \sum_{t=1}^{T}\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}^{m}(x_{t})}$ $(\text{$\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$ and $0\leq\mu_{t}^{m}(x_{t})\leq 1$})$
$\leq$ $\ \sum_{t=1}^{T}\log\left(\frac{\mu_{t}^{m}(x_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $\Big{(}\text{$1-\frac{1}{a}\leq\log a$ for any $a>0$}\Big{)}$
$=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Properties of logarithms})$
$=$ $\ R_{T}^{\times}$ $(\text{Definition of $R_{T}^{\times}$})$
@@
$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
$=$ $\ \sum_{t=1}^{T}\log\left(\frac{\mu_{t}^{m}(x_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{Properties of logarithms})$
$\leq$ $\ \sum_{t=1}^{T}\left(\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{$\log a\leq a-1$ for any $a>0$})$
$\leq$ $\ \sum_{t=1}^{T}\frac{\mu_{t}(x_{t},y_{t})}{\min_{i\in[T]}\mu_{i}(x_{i},y_{i})}\left(\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{$\mu_{t}(x_{t},y_{t})\geq\min_{i\in[T]}\mu_{i}(x_{i},y_{i})>0\ \forall t\in[T]$})$
$=$ $\ \frac{1}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$ $(\text{Arithmetic})$
$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ $(\text{Definition of $R_{T}^{+}$})$
@@
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$. Also, $W/2\leq E[Z]/2$ implies $\Pr[Z\leq W/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
@@
$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ $(\text{Definition of $R_{T}^{+}$})$
@@
$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}\left(1-1+\frac{L}{4f(T)}\right)$ $(\text{bound on $\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ for $t\in M_{j}^{\prime}$})$
$=$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}^{\prime}|}{4f(T)}$ $(\text{Simplifying inner sum})$
@@
$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$ $(\text{$\operatorname*{\mathbb{E}}[Q_{T}]\leq g(T)$})$
$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT}{2^{8}f(T)}\right)$ $(\text{$g(T)\leq f(T)/2$})$
@@
[…]
@@
[…]
@@
  for $t$ from 11 to $T$ do
   with probability $p:$ $\texttt{hedgeQuery}\leftarrow\texttt{true}$
   with probability 1p:1-p: $\texttt{hedgeQuery}\leftarrow\texttt{false}$
@@ B.2 Main proof
[…]
@@
[…]
@@
Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$. Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $a-b||>\delta$. The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
@@
Next, fix a $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the multiset of observed inputs whose mentor action is $y$. Also let $\hat{X}_{2}=\{x_{t}:t\in\hat{Q}_{2}\}$ be the multiset of inputs associated with time steps in $\hat{Q}_{2}$. Note that $|\hat{X}_{2}|=|\hat{Q}_{2}|$, since $\hat{X}_{2}$ is a multiset. We claim that $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$. Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $x-x^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried. Since $x^{\prime}\in\hat{X}_{2}$, this implies $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$. Also, since $x,x^{\prime}\in\hat{X}_{2}$ we have $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$. Therefore
[…]
which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$.
By Lemma B.5, there exists a ball $B_{1}$ of radius $R:=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also an $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.4 implies that
@@
(The +1+1 is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound. Therefore
@@
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
$\ \leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$